A monopolist has a total cost function of 4x2, and faces demand of x = 24 - p. In this case, the optimal output x* for the monopolist to choose is?

Group of answer choices

A. 2.4

B. 4.4

C. 6

D. 8.4
Answer #1

for optional output :- (MCZ MR Here Tc² yuz Mo- dITU du ТМ 3 ) Now, x = 24- P. TP = 24-u] Now, . Total Revenue, ITR = Pxus -